OpenStudy (anonymous):

Find the surface area of the portion of the sphere x^2+y^2+z^2=4 which lies above the xy-plane and BELOW the plane z=1. I just need help with the radius part, please! (i'm using polar coordinates)

OpenStudy (anonymous):

Can you help please?

OpenStudy (anonymous):

You have to use polar coordinates?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Is radius from 0 to 1? and theta from 0to 2pi?

OpenStudy (anonymous):

No, project the part of the sphere over the plane z=1 onto the xy-plane => x^2+y^2=4-z^2=3. r is from -sqrt(3) to sqrt(3), your theta is correct.

OpenStudy (anonymous):

polar coordinates... I figured because it's surface area, It uses double integrals.

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

wait.. why wouldn't it be 1 to sqrt3 because z =1? or z=1 is just used to plug it in the given equation to find out +-sqrt3

OpenStudy (anonymous):

oops, sorry, r would be from 0 to sqrt(3), my bad

OpenStudy (anonymous):

oh, as for your question about the limits of integration, since the projection is a circle of radius sqrt(3) on the xy-plane, we integrate from r=0 to r=sqrt(3). The limits of integration you said is correct if the projection is annular with inner radius=1 and outer radius=sqrt(3).

OpenStudy (anonymous):

OKay... So r is from 1 to sqrt3?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

you just said my limits of integration are correct

OpenStudy (anonymous):

'The limits of integration you said is correct IF the projection is annular with inner radius=1 and outer radius=sqrt(3). (In this case it's not.)

OpenStudy (anonymous):

Oh alright... Thank you!